OpenStudy (anonymous):

Find the limit: lim as x approaches 0 of cos(1/x)

OpenStudy (anonymous):

DNE

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\lim_{x \rightarrow ~0}\cos\left(\frac{1}{x}\right)}\)

OpenStudy (solomonzelman):

That, does not exist

OpenStudy (anonymous):

@pgpilot326 can you explain?

OpenStudy (anonymous):

@solomonzelman why?

OpenStudy (zale101):

If you sub in x=0, you'll get a 0 in the denominator which is undefined.

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\lim_{x \rightarrow ~0}\cos\left(\frac{1}{x}\right)}=\cos\left(\displaystyle \lim_{x \rightarrow ~0}\frac{1}{x}\right)\)

OpenStudy (anonymous):

as x approaches 0, 1/x approaches infinity. cos will cycles through all of it's values and not settle on a single value (which it would need to do in order for the limit to exist)

OpenStudy (solomonzelman):

yes, that is equivalent of \(\large\color{slate}{\displaystyle\lim_{x \rightarrow ~0}\cos\left(\frac{1}{x}\right)}=\cos\left(\displaystyle \lim_{x \rightarrow ~0}\frac{1}{x}\right)=\cos\left(\displaystyle \lim_{x \rightarrow ~\infty }x\right)\)

OpenStudy (solomonzelman):

So it will alternate between 1 and -1

OpenStudy (solomonzelman):